OpenStudy (anonymous):

So You are flipping 2 coins at the same time for 100 times. The results you can get are two head, two tails or one head and one tail. But what does it mean when their asking about frequency? like 27 out of 100 coin flips??

OpenStudy (amistre64):

frequency, how many times do you expect something to occur ...

OpenStudy (amistre64):

hh ht th tt hh is expected 1/4 of the time: freq = 100(1/4) tt is expected 1/4 of the time: freq = 100(1/4) th is expected 1/2 the time: freq = 100(1/2)

OpenStudy (amistre64):

the sum of the freqs is the total of number of events

OpenStudy (amistre64):

if its an experimental run, then its just the number of times it occured during the experiment

OpenStudy (anonymous):

Well I'm supposed to flip 2 coins (at the same time) 100 times and then record my answers.

OpenStudy (amistre64):

start flippin :)

OpenStudy (anonymous):

And how would you know that hh is expected those times. I already flipped them

OpenStudy (amistre64):

theoretical probabilty ..

OpenStudy (amistre64):

the only possible outcomes are: hh th ht tt hh has 1 out of 4 chances ...

OpenStudy (anonymous):

the theoretical probability would be 1/3 because there are 3 results you could get: Two heads, two tails, and one head and one tail

OpenStudy (amistre64):

no, thats not how you work it

OpenStudy (amistre64):

th and ht are both possible outcomes with the same effect

OpenStudy (amistre64):

th and ht happen 1/2 the time in theory

OpenStudy (anonymous):

then how? Can u try to explain it with my problem

OpenStudy (amistre64):

your problem is an experiement, there is no explaining it other than counting up the results.

OpenStudy (amistre64):

we would expect, as you do 1000s and 1000s of flips, that the experimental probability would approach the theoretical probability ...

OpenStudy (anonymous):

Well I'm trying to figure out the theoretical and experimental probability of events while doing the experiment. I'm saying that the theoretical probability would be 1/3 because there are 3 results you could get: Two heads, two tails, and one head and one tail. But idk the experimental probability. I also don't know what it's talking about when it's say's frequency.

OpenStudy (amistre64):

frequency, counting the number of times something occurs ... lets make a table ....